# Neuralized Fermionic Tensor Networks for Quantum Many-Body Systems

### v2

@@ Introduction
In the current work, we are concerned with describing fermionic quantum many-body problems which have a sign structure in the wavefunction. Fermionic TNS [10, 11, 12, 13, 14, 15, 16, 17, 18] utilize a $\mathbb{Z}_{2}$ graded tensor product to encode the sign structure and have been demonstrated to be a powerful Ansatz for many problems [19, 15, 20, 21, 22, 16, 23]. At the same time, TNFs allow for substantial flexibility in function design, including the incorporation of nonlinearity into the tensor network computational graph. Motivated by both these observations, we propose a class of fermionic TNFs enhanced with neural network-induced nonlinearity, which we denote NN-fTNS. NN-fTNS serve as a bridge between pure tensor networks and pure neural networks, combining the conceptual advantages of tensor networks with the expressive flexibility of neural networks. While hybrid Ansätze [24, 25, 26, 27, 28, 29] combining tensor networks and neural networks have been explored in bosonic spin systems [25, 26, 29], here we demonstrate the power of this hybrid approach in the fermionic setting, in the framework of tensor network functions, where the tensor network arguably plays a more critical role. In particular, in the study of fermionic neural network quantum states (NQS) [4, 5, 30, 31, 32], it is essential to incorporate a proper physical bias towards the correct fermionic sign structure. Most conventional fermionic NQS approaches rely on a mean-field Slater determinant or Pfaffian to provide this bias [5, 31, 32, 33, 30]. NN-fTNS use the fTNS computational graph arising from the $\mathbb{Z}_{2}$ graded tensor product as the physical bias for the neural network function, providing a conceptually distinct option from existing mean-field based NQS.
@@
Figure 2: VMC optimization curve of various relevant Ansätze for the FH model at hole doping $n_{h}=1/4$ on a small $4\times 2$ square lattice. Errors are calculated relative to the exact ground state energy. NN: pure neural network parametrization; NN-SD: optimized Slater determinant with NN backflow; fPEPS: pure fPEPS state; NN-fPEPS: neuralized fPEPS; fPEPS×$\times$×NN-Jastrow: Multiplicative NN-Jastrow factor on fPEPS.
@@
To first demonstrate that fTNS provide a suitable fermionic sign structure, we compare the ground state optimization performance of NN-fTNS against various other Ansatz on a small $4\times 2$ FH model on square lattice with hole doping $n_{h}=1/4$; see the VMC optimization curves in Fig. 2, details of all Ansatz in the SM [42]; the NN model sizes (which dominate the number of parameters) for all Ansatz were chosen to be similar, and the fPEPS bond dimension is fixed to $D=4$ in all fTNS-based models. Comparing NN-fPEPS with pure fPEPS (initialization obtained by SU), we clearly observe that the addition of the neural network configuration dependence increases the model expressivity, allowing a further energy reduction beyond what is achieved by pure fPEPS. Comparing NN-fPEPS with a pure randomly initialized NN Ansatz, it is evident that the fTNS provides the crucial fermionic sign structure that the pure neural network fails to capture, with the latter yielding a large relative energy error of approximately $30\%$ despite a similar model size. For a further comparison, we also use a conventional neural network backflow state [5], where a Slater determinant is used to encode the fermionic sign structure (denoted as NN-SD). We note that this architecture is similar to that in Ref. [5] and is not intended to represent the state-of-the-art in NN backflow [30, 33, 41, 52], but provides a baseline NQS calculation. Starting from a fully optimized Slater determinant and subsequently adding and optimizing the neural network backflow, our NN-SD calculation achieves a relative error of only about $1\%$. In contrast, the NN-fPEPS model reaches a relative error below $5\times 10^{-4}$ in fewer than 400 optimization steps, illustrating the power of fTNS as a fermionic bias. To verify that the NN modulation of this bias is important, we have also tested adding a neural network to the fPEPS through a multiplicative Jastrow factor [25, 53]. While this does improve over the pure fPEPS calculation, the improvement is quite limited, showing the importance of neural network modulation of the sign structure.
@@
In contrast, removing the fTNS bias results in significantly worse energies. A pure neural network Ansatz yields relative energy errors of 22.68% and 14.78% for the $4\times 4$ FH model at half filling and at hole doping $n_{h}=1/8$, respectively [42], and a recent study using an autoregressive NQS without a fermionic mean-field bias for the same system at half filling reported a relative error of $8\times 10^{-3}$ [55], worse than our smallest NN-fPEPS calculation with bond dimension $D=4$, which achieved a $4\times 10^{-4}$ relative error. Similarly, a baseline Slater determinant neural network backflow calculation achieved relative energy errors of only 2.31% at half filling and 3.18% at $n_{h}=1/8$ [42]. These results highlight the quality of the fermionic sign structure provided by the fTNS.
Figure 3: Relative energy error w.r.t. DMRG energy of the FH model for different fTNS bond dimension $D$ with various lattice geometries and hole dopings. (a) $24$-site chain at hole doping $n_{h}=1/6$. (b) $4\times 4$ square lattice at half filling. (c) $4\times 4$ square lattice at $n_{h}=1/8$ hole doping. (d) $6\times 6$ square lattice at $n_{h}=1/9$ hole doping. DMRG energy for the one-dimensional FH model is obtained with a fMPS of bond dimension $m=1000$. DMRG and fPEPS energies for the two-dimensional FHM are taken from Ref. [19].
@@
The cost to compute an amplitude in a NN-fTNS is $O(N)$ where $N$ is the system size [44, 19, 56]. This can be contrasted with the $O(N^{3})$ scaling for a Slater determinant or Pfaffian NQS [5]. However, in a Markov Chain VMC calculation, typically a series of related configurations are generated in a sweep [57], for example in the computation of the local energy. In a pure fTNS VMC calculation, computation can be reused (in the form of partial tensor network contraction intermediates [42]) which means that for a local Hamiltonian (such as the FH model), the cost per VMC energy and step is still $O(N)$ [58, 6]. However, the global configuration dependence of a NN-fTNS means that even local configuration changes in the amplitude lead to a global modification of the tensors, preventing re-use and leading to an $O(N^{2})$ scaling for each VMC step. (This is related to the $O(N^{4})$ scaling of a VMC step with Slater determinant/Pfaffian NQS when reuse is not possible [33, 52, 32, 30]). This is illustrated in the time cost comparisons between NN-fPEPS and fPEPS for various bond dimensions $D$ in Fig. 5(a). Clearly, there exists a trade-off between the range of neural network-induced configuration dependence and the degree of reuse of partial tensor network contractions.
@@
Figure 5: (a) Time cost comparison for a single VMC step ($20$ samples per step) across different fPEPS bond dimensions $D$ on a $6\times 6$ FH model. Each timing is averaged over 5 independent trials on a single CPU core. The boundary MPS contraction bond dimension is chosen as $\chi=4D$. (b) Scalings of the time cost per VMC step ($20$ samples per step) with system size $N$, using NN-fMPS ($D=6$) models with different neural network range $R$. Each timing is averaged over 5 independent trials on a single CPU core. (c)-(d): Relative ground-state energy errors of NN-fTNS with varying neural network ranges $R$ for the FH model on a (c) $24$-site chain and (d) $6\times 6$ square lattice. In (d), the fPEPS benchmark energy is taken from Ref. [19].
@@ Acknowledgments.
We thank Ao Chen for sharing data on pure NN and NNBF calculations. SD thanks Wen-Yuan Liu, Ruojing Peng, Johnnie Gray, Di Luo and Ao Chen for helpful discussions. The fermionic tensor networks are implemented with Quimb [59] and Symmray [10], and neural networks are implemented with PyTorch [60]. This work was supported by the US Department of Energy, Office of Science, through Award no. DE-SC0019374. GKC acknowledges additional support from the Simons Investigator program.
@@
[…]
@@
The features $X$ are subsequently processed by a multi-head self-attention block with residual connection [39]:
@@
For the NN-SD model, we follow the standard neural network backflow construction [5, 52], replacing the fTNS in Fig. S1(a) with a single-particle orbital matrix $M_{i,k}\in\mathbb{R}^{N_{o}\times N_{e}}$, where $N_{o}$ denotes the total number of spin orbitals and $N_{e}$ is the total number of electrons. Given a configuration $|n\rangle$, the NN modifies the orbital matrix to produce a configuration-dependent matrix $M(n)$. The rows of $M(n)$ corresponding to the occupied orbitals in $|n\rangle$ are selected to form a square matrix, whose determinant $\det(M_{i_{o},j})$ yields the final amplitude, where $i_{o}$ labels the occupied orbitals. In VMC optimization, NN-SD is initialized by first fully optimizing the Slater determinant orbital matrix (also via VMC), after which the NN parameters are initialized such that the neural network introduces only small perturbations to the fully optimized Slater determinant.
For the fPEPS×$\times$×NN-Jastrow model, given a configuration $|n\rangle$ and an fPEPS state $|\Psi\rangle$, we apply the pure NN model to compute a scalar factor $\mathrm{NN}(n)$, which is multiplied with the fPEPS amplitude to obtain the final amplitude $\mathrm{NN}(n)\times\langle n|\Psi\rangle$. In VMC optimization, the fTNS tensors are initialized using SU, and the NN model parameters are initialized randomly.
@@
In VMC, operator expectation values are estimated via importance sampling of configurations $|n\rangle=|n_{1},\dots,n_{N}\rangle$, where $N$ is the system size. The energy of a Hamiltonian $H$ is evaluated as
@@
For fTNS based models where local tensors only have local configuration dependence, partial TN contraction can be reused for computation of related amplitudes, as illustrated in Fig.S2. For pure fTNS, this reuse strategy allows the VMC cost to be reduced from $O(N^{2})$ to nearly $O(N)$ in principle. For example, in the computation of the local energy for a configuration $|n\rangle$ using boundary MPS contraction, one may first cache all partial TN contraction intermediates (environment MPS $|E_{i}\rangle$ in Fig.S2) for $\langle n|\Psi\rangle$. Then, to compute $\langle n_{c}|\Psi\rangle$ for a locally connected configuration $|n_{c}\rangle$, only a three-row TN $\langle E_{i-1}|\hat{O}_{i}|E_{i+1}\rangle$ needs to be contracted, which scales as $O(\sqrt{N})$ for a square lattice. If additional TN reuse along the column direction is employed, each $|n_{c}\rangle$ requires contracting only a small plaquette TN with constant cost $C_{n_{c}}$ independent of system size, reducing the total cost for local energy evaluation to $O(C_{n_{c}}N)$ [44, 58, 6]. For NN-fTNS models with finite neural network range $R$, where each local tensor depends on neighboring configurations within range $R$, a similar partial TN contraction reuse scheme can be applied.
@@
Model D $4$ $6$ $8$
fMPS-SU $-0.490(2)$ $-0.531(7)$ $-0.532(8)$
fMPS-GO $-0.5037(8)$ $-0.5348(0)$ $-0.5470(4)$
@@
Table 1: Energy per site of pure fMPS and NN-fMPS for the Fermi-Hubbard model on an $L=24$ one-dimensional chain at hole doping $n_{h}=1/6$ ($N_{e}=20$, $N_{\uparrow}=N_{\downarrow}=10$). The DMRG reference energy per site is $-0.56309964$, obtained using pyBlock2 [68] with fMPS bond dimension 1000 and truncated weight $\mathrm{d}w=1.6\times 10^{-10}$.
@@
Model D $4$ $6$ $8$ $14$
fPEPS-SU [19] $-0.4174(2)$ $-0.4204(6)$ $-0.4238(5)$ $-0.4242(5)$
fPEPS-GO [19] $-0.42115(3)$ $-0.42416(7)$ $-0.42508(4)$ $-0.425508(5)$
NN-fPEPS-GO $\mathbf{-0.42535\pm 4\times 10^{-5}}$ $\mathbf{-0.42541\pm 3\times 10^{-5}}$ $\mathbf{-0.42551\pm 2\times 10^{-5}}$
Table 2: Energy per site of pure fPEPS and NN-fPEPS for the Fermi-Hubbard model on a $4\times 4$ square lattice at half-filling ($N_{e}=16$, $N_{\uparrow}=N_{\downarrow}=8$). The DMRG reference energy per site is $-0.42552590$ [19]. The pure NN model yields $E_{\mathrm{NN}}=-0.3290$, while the NN backflow baseline calculation gives $E_{\mathrm{NN-SD}}=-0.4157$.
@@
Model D $4$ $6$ $8$ $10$ $14$
fPEPS-SU [19] $-0.5497(6)$ $-0.5850(4)$ $-0.6016(3)$ $-0.6104(4)$ $-0.6154(4)$
fPEPS-GO [19] $-0.5892(2)$ $-0.61225(7)$ $-0.62314(6)$ $-0.62725(4)$ $-0.63206(1)$
NN-fPEPS-GO $\mathbf{-0.6247\pm 2\times 10^{-4}}$ $\mathbf{-0.6301\pm 2\times 10^{-4}}$ $\mathbf{-0.63215\pm 7\times 10^{-5}}$ $\mathbf{-0.63231\pm 5\times 10^{-5}}$
Table 3: Energy per site of pure fPEPS and NN-fPEPS for the Fermi-Hubbard model on a $4\times 4$ square lattice at $1/8$ hole doping ($N_{e}=14$, $N_{\uparrow}=N_{\downarrow}=7$). The DMRG reference energy is $-0.63261830$ [19]. The pure NN model yields $E_{\mathrm{NN}}=-0.5391$, while the NN backflow baseline calculation gives $E_{\mathrm{NN-SD}}=-0.6125$.
@@
Model D $4$ $6$ $8$ $16$
fPEPS-SU[19] $-0.5786(4)$ $-0.6200(6)$ $-0.6312(5)$ $-0.6470(4)$
fPEPS-GO[19] $-0.6007(7)$ $-0.6309(4)$ $-0.6415(6)$ $-0.65327(5)$
NN-fPEPS-GO $\mathbf{-0.6370\pm 2\times 10^{-4}}$ $\mathbf{-0.6503\pm 1\times 10^{-4}}$ $\mathbf{-0.6535\pm 1\times 10^{-4}}$
Table 4: Energy per site of pure fPEPS and NN-fPEPS for the Fermi-Hubbard model on a $6\times 6$ square lattice at $1/9$ hole doping ($N_{e}=32$, $N_{\uparrow}=N_{\downarrow}=16$). The DMRG reference energy per site is $-0.6573063$ [19].

### v3

@@ Introduction
In the current work, we are concerned with describing fermionic quantum many-body problems which have a sign structure in the wavefunction. Fermionic TNS [10, 11, 12, 13, 14, 15, 16, 17, 18] utilize a $\mathbb{Z}_{2}$ graded tensor product to encode the sign structure and have been demonstrated to be a powerful Ansatz for many problems [19, 15, 20, 21, 22, 16, 23]. At the same time, TNFs allow for substantial flexibility in function design, including the incorporation of nonlinearity into the tensor network computational graph. Motivated by both these observations, we propose a class of fermionic TNFs enhanced with neural network-induced nonlinearity, which we denote NN-fTNS. NN-fTNS serve as a bridge between pure tensor networks and pure neural networks, combining the conceptual advantages of fermionic tensor networks with the expressive flexibility of neural networks. While hybrid Ansätze [24, 25, 26, 27, 28, 29] combining tensor networks and neural networks have been explored in bosonic spin systems [25, 26, 29], here we demonstrate the power of this hybrid approach in the fermionic setting, in the framework of tensor network functions, where the tensor network arguably plays a more critical role. In particular, in the study of fermionic neural network quantum states (NQS) [4, 5, 30, 31, 32], it is essential to incorporate a proper physical bias towards the correct fermionic sign structure. Most conventional fermionic NQS approaches rely on a mean-field Slater determinant or Pfaffian to provide this bias [5, 31, 32, 33, 30]. NN-fTNS use the fTNS computational graph arising from the $\mathbb{Z}_{2}$ graded tensor product as the physical bias for the neural network function, providing a conceptually distinct option from existing mean-field based NQS.
@@
[…]
@@
To first demonstrate that fTNS provide a suitable fermionic sign structure, we compare the ground state optimization performance of NN-fTNS against various other Ansätze on a small $4\times 2$ FH square lattice with hole doping $n_{h}=1/4$; the VMC converged energy is shown as a function of the number of variational parameters in Fig. 2, and details of all Ansätze in the SM [42]; the fPEPS bond dimension is fixed to $D=4$ in all fTNS-based models. The NN model sizes dominate the number of parameters for all Ansätze. Comparing NN-fPEPS with pure fPEPS (initialized by SU), we clearly observe that the addition of the neural network configuration dependence increases the model expressivity, allowing a further energy reduction beyond what is achieved by pure fPEPS. The pure NN Ansatz reaches a large relative energy error of approximately $30\%$ despite a large model size, demonstrating that the fTNS provides the crucial fermionic sign structure that the pure neural network fails to capture. For a further comparison, we also use a conventional neural network backflow state [5], where a Slater determinant is used to encode the fermionic sign structure (denoted as NN-SD). We implement this NN-SD calculation with two NN architectures: (1) a simple feed-forward network (NN-SD (MLP)), similar to Ref. [5], and (2) a modern Transformer [39, 52] (NN-SD (ViT)), which mimics the sophisticated NQS model structure in recent state-of-the-art work [53]. As shown in Fig. 2, while increasing the NN size reduces the NN-SD energy, the NN-fPEPS model achieves a lower relative error for a comparable number of variational parameters. Additionally, the NN-fPEPS model requires fewer VMC optimization steps to converge than the NN-SD model (see SM [42] for VMC optimization curves). To verify that the NN modulation of the fermionic bias from fTNS is important, we have also tested adding a neural network to the fPEPS through a multiplicative Jastrow factor [25, 54]. While this does improve over the pure fPEPS calculation, the improvement is quite limited, showing the importance of direct neural network modulation of the sign structure.
@@
In contrast, removing the fTNS bias results in significantly worse energies. A pure neural network Ansatz yields relative energy errors of 22.68% and 14.78% for the $4\times 4$ FH model at half filling and at hole doping $n_{h}=1/8$, respectively [42], and a recent study using an autoregressive NQS without a fermionic mean-field bias for the same system at half filling reported a relative error of $8\times 10^{-3}$ [56], worse than our smallest NN-fPEPS calculation with bond dimension $D=4$, which achieved a $4\times 10^{-4}$ relative error. Similarly, a baseline NN-SD calculation achieved relative energy errors of only 2.31% (MLP), 0.18% (ViT) at half filling and 3.18% (MLP), 0.6% (CNN [42, 57]) at $n_{h}=1/8$ [42]. These results highlight the quality of the fermionic sign structure provided by the fTNS.
Figure 3: Relative energy error w.r.t. DMRG energy of the FH model for different fTNS bond dimension $D$ with various lattice geometries and hole dopings. (a) $24$-site chain at hole doping $n_{h}=1/6$. (b) $4\times 4$ square lattice at half filling. (c) $4\times 4$ square lattice at $n_{h}=1/8$ hole doping. (d) $6\times 6$ square lattice at $n_{h}=1/9$ hole doping. DMRG energy for the one-dimensional FH model is obtained with a fMPS of bond dimension $m=1000$. DMRG and fPEPS energies for the two-dimensional FHM are taken from Ref. [19]. In (b) and (c), we provide the baseline energies of NN-SD calculations with sophisticated model architectures for comparison. The NN-SD baseline models each have about 100,000 parameters, a size comparable to that of the NN-fPEPS model with $D=6$.
@@
The cost to compute an amplitude in a NN-fTNS is $O(N)$ where NN is the system size [44, 19, 58]. This can be contrasted with the $O(N^{3})$ scaling for a Slater determinant or Pfaffian NQS [5]. However, in a Markov Chain VMC calculation, typically a series of related configurations are generated in a sweep [59], for example in the computation of the local energy. In a pure fTNS VMC calculation, computation can be reused (in the form of partial tensor network contraction intermediates [42]) which means that for a local Hamiltonian (such as the FH model), the cost per VMC energy and step is still $O(N)$ [60, 6]. However, the global configuration dependence of a NN-fTNS means that even local configuration changes in the amplitude lead to a global modification of the tensors, preventing re-use and leading to an $O(N^{2})$ scaling for each VMC step. (This is related to the $O(N^{4})$ scaling of a VMC step with Slater determinant/Pfaffian NQS when reuse is not possible [33, 61, 32, 30]). This is illustrated in the time cost comparisons between NN-fPEPS and fPEPS for various bond dimensions $D$ in Fig. 5(a). Clearly, there exists a trade-off between the range of neural network-induced configuration dependence and the degree of reuse of partial tensor network contractions.
@@
Figure 5: (a) Time cost comparison for a single VMC step (2020 samples per step) across different fPEPS bond dimensions $D$ on a $6\times 6$ FH model. Each timing is averaged over 5 independent trials on a single CPU core. The boundary MPS contraction bond dimension is chosen as $\chi=4D$. (b) Scalings of the time cost per VMC step (2020 samples per step) with system size NN, using NN-fMPS ($D=6$) models with different neural network range $R$. Each timing is averaged over 5 independent trials on a single CPU core. (c)-(d): Relative ground-state energy errors of NN-fTNS with varying neural network ranges $R$ for the FH model on a (c) $24$-site chain and (d) $6\times 6$ square lattice. In (d), the fPEPS benchmark energy is taken from Ref. [19].
@@ Acknowledgments.
SD thanks Wen-Yuan Liu, Ruojing Peng, Johnnie Gray and Di Luo for helpful discussions. The fermionic tensor networks are implemented with Quimb [62] and Symmray [10], and neural networks are implemented with PyTorch [63]. This work was supported by the US Department of Energy, Office of Science, through Award no. DE-SC0019374. Some computations were performed using the facilities of National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, under NERSC award ERCAP0023924. GKC acknowledges additional support from the Simons Investigator program.
@@
NN-fTNS computational graph We describe the computational graph shown in Fig. S1(a) from bottom to top. The local on-site configuration of the $i$-th site is first encoded using a one-hot representation, where the four basis states $\{0,\uparrow,\downarrow,\uparrow\downarrow\}$ are mapped to the four orthonormal unit vectors $\mathbf{e}_{i}^{(j)}\in\mathbb{R}^{4}$ ($j=0,1,2,3$) in four-dimensional Euclidean space. The one-hot encoded on-site configuration is then embedded into a $d$-dimensional feature vector (Fig.S1(b)) through a linear transformation:
@@
The features $X$ are subsequently fed into a standard Transformer encoder, which has an analogous structure to a Vision Transformer (ViT) [52]. Specifically, $X$ are input into a multi-head self-attention block with residual connection [39]:
@@
For the NN-SD model, we follow the standard neural network backflow construction [5, 61, 53], replacing the fTNS in Fig. S1 with a single-particle orbital matrix $M_{i,k}\in\mathbb{R}^{N_{o}\times N_{e}}$, where $N_{o}$ denotes the total number of spin orbitals and $N_{e}$ is the total number of electrons. Given a configuration $|n\rangle$, the NN modifies the orbital matrix to produce a configuration-dependent matrix $M(n)$. The rows of $M(n)$ corresponding to the occupied orbitals in $|n\rangle$ are selected to form a square matrix, whose determinant $\det(M_{i_{o},j})$ yields the final amplitude, where $i_{o}$ labels the occupied orbitals. In VMC optimization, the NN-SD is initialized by first fully optimizing the Slater determinant orbital matrix (via minimizing its mean-field energy in the FH model), after which the NN parameters are initialized such that the neural network introduces only small perturbations to the fully optimized Slater determinant. The specific NN structures used are (1) NN-SD (MLP): a two-layer multilayer perceptron; (2) NN-SD (ViT): a Vision Transformer [52] as shown in Fig. S1(a); (3) a convolutional neural network [57, 65].
[…]
S-1.3 VMC optimization curves for various Ansätze
We show below an example of our VMC optimization procedure for various Ansätze on a small $4\times 2$ FH model. We can see the NN-fPEPS model requires fewer VMC optimization steps (around 500 steps) to converge than the NN-SD model (at least 1000 steps).
[…]
@@
In VMC, operator expectation values are estimated via importance sampling of configurations $|n\rangle=|n_{1},\dots,n_{N}\rangle$, where NN is the system size. The energy of a Hamiltonian $H$ is evaluated as
@@
For fTNS based models where local tensors only have local configuration dependence, partial TN contraction can be reused for computation of related amplitudes, as illustrated in Fig.S3. For pure fTNS, this reuse strategy allows the VMC cost to be reduced from $O(N^{2})$ to nearly $O(N)$ in principle. For example, in the computation of the local energy for a configuration $|n\rangle$ using boundary MPS contraction, one may first cache all partial TN contraction intermediates (environment MPS $|E_{i}\rangle$ in Fig.S3) for $\langle n|\Psi\rangle$. Then, to compute $\langle n_{c}|\Psi\rangle$ for a locally connected configuration $|n_{c}\rangle$, only a three-row TN $\langle E_{i-1}|\hat{O}_{i}|E_{i+1}\rangle$ needs to be contracted, which scales as $O(\sqrt{N})$ for a square lattice. If additional TN reuse along the column direction is employed, each $|n_{c}\rangle$ requires contracting only a small plaquette TN with constant cost $C_{n_{c}}$ independent of system size, reducing the total cost for local energy evaluation to $O(C_{n_{c}}N)$ [44, 60, 6]. For NN-fTNS models with finite neural network range $R$, where each local tensor depends on neighboring configurations within range $R$, a similar partial TN contraction reuse scheme can be applied. A comparison of the VMC cost scaling for the above mentioned models as well as conventional fermionic NQS is listed in Table. 1.
@@
Table 1: Comparison of VMC cost scaling with respect to system size NN for various fermionic Ansätze.
Model VMC cost scaling
NN-fTNS (no reuse) $\mathcal{O}(N^{2})$
NN-fTNS (with reuse) $\geq O(N),\ <\mathcal{O}(N^{2})$
fTNS (with reuse) $\mathcal{O}(N)$
NN-SD (Ref. [5]) $\mathcal{O}(N^{4})$
HFPS (Ref. [57]) $\mathcal{O}(N^{3})$
@@
Model D 44 66 88
fMPS-SU 0.490(2)-0.490(2) 0.531(7)-0.531(7) 0.532(8)-0.532(8)
fMPS-GO 0.5037(8)-0.5037(8) 0.5348(0)-0.5348(0) 0.5470(4)-0.5470(4)
@@
Table 2: Energy per site of pure fMPS and NN-fMPS for the Fermi-Hubbard model on an $L=24$ one-dimensional chain at hole doping $n_{h}=1/6$ ($_{e}=20, N=N$$_{\uparrow}=N_{\downarrow}=10$). The DMRG reference energy per site is 0.56309964-0.56309964, obtained using pyBlock2 [72] with fMPS bond dimension 1000 and truncated weight $\mathrm{d}w=1.6\times 10^{-10}$.
@@
Model D 44 66 88 $14$
fPEPS-SU [19] 0.4174(2)-0.4174(2) 0.4204(6)-0.4204(6) 0.4238(5)-0.4238(5) 0.4242(5)-0.4242(5)
fPEPS-GO [19] 0.42115(3)-0.42115(3) 0.42416(7)-0.42416(7) 0.42508(4)-0.42508(4) 0.425508(5)-0.425508(5)
NN-fPEPS-GO $\mathbf{-0.42535\pm 4\times 10^{-5}}$ $\mathbf{-0.42541\pm 3\times 10^{-5}}$ $\mathbf{-0.42551\pm 2\times 10^{-5}}$ -
Table 3: Energy per site of pure fPEPS and NN-fPEPS for the Fermi-Hubbard model on a $4\times 4$ square lattice at half-filling ($_{e}=16, N=N$$_{\uparrow}=N_{\downarrow}=8$). The DMRG reference energy per site is 0.42552590-0.42552590 [19]. The pure NN model yields $E_{\mathrm{NN}}=-0.3290$, while the NN backflow baseline calculation gives $E_{\mathrm{NN-SD\ (MLP)}}=-0.4157$ and $E_{\mathrm{NN-SD\ (ViT)}}=-0.4247$.
@@
Model D 44 66 88 $10$ $14$
fPEPS-SU [19] 0.5497(6)-0.5497(6) 0.5850(4)-0.5850(4) 0.6016(3)-0.6016(3) 0.6104(4)-0.6104(4) 0.6154(4)-0.6154(4)
fPEPS-GO [19] 0.5892(2)-0.5892(2) 0.61225(7)-0.61225(7) 0.62314(6)-0.62314(6) 0.62725(4)-0.62725(4) 0.63206(1)-0.63206(1)
NN-fPEPS-GO $\mathbf{-0.6247\pm 2\times 10^{-4}}$ $\mathbf{-0.6301\pm 2\times 10^{-4}}$ $\mathbf{-0.63215\pm 7\times 10^{-5}}$ $\mathbf{-0.63231\pm 5\times 10^{-5}}$ -
Table 4: Energy per site of pure fPEPS and NN-fPEPS for the Fermi-Hubbard model on a $4\times 4$ square lattice at $1/8$ hole doping ($_{e}=14, N=N$$_{\uparrow}=N_{\downarrow}=7$). The DMRG reference energy is 0.63261830-0.63261830 [19]. The pure NN model yields $E_{\mathrm{NN}}=-0.5391$, while the NN backflow baseline calculation gives $E_{\mathrm{NN-SD\ (MLP)}}=-0.6125$ and $E_{\mathrm{NN-SD\ (CNN)}}=-0.6288$.
@@
Model D 44 66 88 $16$
fPEPS-SU[19] 0.5786(4)-0.5786(4) 0.6200(6)-0.6200(6) 0.6312(5)-0.6312(5) 0.6470(4)-0.6470(4)
fPEPS-GO[19] 0.6007(7)-0.6007(7) 0.6309(4)-0.6309(4) 0.6415(6)-0.6415(6) 0.65327(5)-0.65327(5)
NN-fPEPS-GO $\mathbf{-0.6370\pm 2\times 10^{-4}}$ $\mathbf{-0.6503\pm 1\times 10^{-4}}$ $\mathbf{-0.6535\pm 1\times 10^{-4}}$ -
Table 5: Energy per site of pure fPEPS and NN-fPEPS for the Fermi-Hubbard model on a $6\times 6$ square lattice at $1/9$ hole doping ($_{e}=32, N=N$$_{\uparrow}=N_{\downarrow}=16$). The DMRG reference energy per site is 0.6573063-0.6573063 [19].